OpenStudy (anonymous):

Simplify completely x2+4x-5 . 20x-12 -------- ------- 5x2-8x+3 x2-6x-55

OpenStudy (callisto):

1. factor the expression. (i) x^2+4x-5 (ii) 5x^2-8x+3 (iii) 20x-12 (iv) x^2-6x-55 What do you get for factoring the above expressions?

OpenStudy (anonymous):

I just need the answer pleezz! :(

OpenStudy (callisto):

@cassie4math I'm sorry that I can't give you the direct answer, but I don't mind guiding you to get your answer!

OpenStudy (callisto):

Hint: For (i) x^2+4x-5 = x^2 -x +5x - 5 = x(x-1) + 5(x-1) = (x+5)(x-1) Can you do the other three?

OpenStudy (anonymous):

is the answer over all 4/x-11.... this is what i got before and i was just checking it on here!

OpenStudy (callisto):

That's what I got :|

OpenStudy (anonymous):

k thank you!

OpenStudy (callisto):

Welcome.